40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 264 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  3806   Thu Oct 28 04:23:38 2010 ranaUpdateIOOA2L prep

To get the angle to length signal before the c1ioo processor gets going, we need a length signal. We can use either the error signal or the control signal.

I recommend using the control signal since its not puny. The 4-pin LEMO inputs to the OSEM ADC that Suresh has wired are differential so we can, in principle, use either the BNC output of the SERVO plug or the 2-pin LEMO output.

The analog whitening on the OSEM Whitening board should be engaged via the SUS MEDM screen so that we get a good SNR at the A2L dither frequencies.

If the ADC saturates, then we should use a pomona box RC low pass to cut everything off above 100 Hz.

Also, a comment about Yuta's elog: we estimated that the seismic motion was ~1e-7 - 1e-6 meters. The MC linewidth ought to be ~lambda/(2*Finesse) ~ 1e-9.

So, the MC servo as it was was not giving us enough gain (1/f above 50 Hz; UGF ~5-10 kHz) to get the error signal to stay in the linear PDH region. Kevin's filter gave us ~10x more gain at the seismic frequencies (1-3 Hz) of concern.

  3805   Thu Oct 28 03:39:58 2010 yutaUpdateIOOgot very stable MC locking

(Rana, Suresh, Jenne, Kiwamu, Kevin, Yuta)

Summary:
  Last night we locked MC by feeding back the signal to NPRO PZT.
  But it was not so stable.
  We wanted more gain to lower the seismic motion, but we don't need high gain in high frequency part(>~1kHz) because it may cause something bad(NRPO PZT oscilatting, PMC not able to catch up with the NPRO frequency change, etc).
  So, we put DC gain boost today.
  It successfully made MC locking stable!

What we did:
1. Lowered the main laser temperature from 32.2° C to 31.8° C.
  When we increased the laser temperature, PMC transmission get lower.  32.2° C was on the cliff, so we put it to plateau region.

2. Lowered the gain of PMC servo (2dB instead of 8dB last night), because PMC was oscillating.
  We got 5.3V OMC transmission.

3. Made 3Hz pole, 30Hz zero filter and put in NPRO PZT servo loop.
  0dB at DC, -20dB at high frequency (see Kevin's elog #3802)

4. Put more gain to NPRO PZT servo loop to compensate -20dB at high frequency.
  In a word, we put DC gain boost.
  Attachment #1 is the MEDM screen screenshot for MC servo.

5. Aligned the beam into QPD at MC2 trans.
  We put lens in front of the QPD.
  Now, we can see the actual motion of the beam, and resonance peaks(Attachment #2; not locked at the highest).
    (We added 30Hz LPF after each 4 quadrant inputs to reduce noise)

Plan:
 - optimize MC suspension alignments
 - activate OL DAQ channels
 - reduce RFAM
 - install tri-mod box
 - QPD signal at MC2 should be more high(currently, ~7counts which equals to ~4mV)
 - change temperature of X-end laser to get green beat

Attachment 1: C1IOO_MC_SERVO20101028.png
C1IOO_MC_SERVO20101028.png
Attachment 2: MCrocks!.png
MCrocks!.png
  3804   Thu Oct 28 03:21:35 2010 SureshUpdateLockingAligned the MC2 transmission photodiode

Yuta and Suresh

The MC2 transmission is seen on the QPD

Once the laser was locked to the cavity, and the PMC was able to follow the laser (ref: elogs by Yuta and Rana today)  we had a steady TEMoo mode in the MC.  This gave us sufficient transmission through MC2 to be easily visible with an IR viewer and we were able to guide the beam on to the QPD.  The beam however seemed to over fill the QPD, a lens (f=180mm) was placed before the beam folding mirror and the spot sized reduced.   The the QPD was found to be not fixed to the table and this was also recitified.  The signal level is still low: total signal is just about 7 DAQ steps amounting to about 5mV.  Tomorrow we plan to work on the alignment of the PSL and MC and thus increase this signal.

A new channel to observe the length variations in the MC.

A long BNC cable was laid from the MC locking electronics next (southwards) to the PSL table to the DAQ cards picking up the signals from the PRM OSEMS.  This is to pick up one of the MC locking signals and collect it on a DAQ channel.  However as there are no spare DAQ channels currently available one of the PRM OSEM (UL and LL) photodiode channels was unplugged and replaced with the signal from the long BNC cable.   For identifying the correct DAQ channel we put in a 2 Vpp 10Hz signal with a function generator into this BNC.  Tow signals can be picked up in this fashion and they are available on PRM_LLSEN_IN1 and PRM_ULSEN_IN1. We plan to use this for improving the alignment of the MC tomorrow. 

The algorithm for this alignment is that if the beam from the PSL is not centered on the MC1 then tilting MC1 would result in a change in the length of the cavity as seen by the light.  Using this as feedback the spot could be precisely centered on the MC1 and then the MC alignment could be completed by moving MC2 and MC3 to reobtain TEM_oo within the cavity.

  3803   Thu Oct 28 03:07:53 2010 kiwamuUpdateGreen Locking80MHz VCO for green PLL : a health check

 I did a health check for a 80MHz VCO box. 

I started taking care with the black VCO box, which has been sitting on the SP table and will be used for converting the green beat signal from frequency to voltage.

The circuit in the box basically consists of three parts: low pass filters (LPFs), a VCO and RF amplifiers.

Today I checked the LPF stage. It looks pretty healthy.

Tomorrow I will check the VCO part, especially I am curious about the VCO range.

 


 (soldering)

 Since somebody ( surf students ?) removed some resistors, the VCO was just freely running without being applied any voltage.

I put some resistors back on the circuit board by soldering them.

Now the resistors are placed in the same configuration as the original schematic (link to LIGO DCC) except for the wideband signal path, which has a differential input.

I left the wideband path disconnected from the VCO.

 

(transfer function measurement)

The LPF part in 'external mod' path contains two stages in series:

one is for cutting off demodulated signals above fc=80MHz and the other one is for PLL servo with pole=1Hz, zero=40Hz.

In order to activate this path I shorted 10th pin of the analog switch: MAX333A.

During the transfer function measurement I injected signals to 'external mod' input and took the output signal from a test point pin TP7.

The plot below shows a fitting result of the measured transfer function of the whole LPF stage. I used liso for the fitting.

The measured filter's shape agreed with the design. (though I haven't checked 80MHz cut off)

VCO_LPF_fit.png

  3802   Thu Oct 28 02:01:51 2010 KevinUpdatePSLFilter for 2W Laser

[Rana and Kevin]

I made a low pass filter for the piezo driver for the 2W laser that is now installed. The filter has a pole at 2.9 Hz. The transfer function is shown in attachment 1.

Attachment 2 shows the outside of the filter with the circuit diagram and attachment 2 shows the inside of the filter.

Attachment 1: tf.PDF
tf.PDF
Attachment 2: outside.jpg
outside.jpg
Attachment 3: inside.jpg
inside.jpg
  3801   Wed Oct 27 21:51:29 2010 SureshUpdatePSLAttenuation of PSL NPRO removed

The laser power was attenuated to 40 mW yesterday for ensuring that the power built up within the MC does not damage the optics. 

This however stopped us from the doubling work and besides also reduced the power available for locking the PMC. 

Therefore, today the laser attenuation was removed and once again 500mW is available at the exit of the PMC .  

However the power sent to the MC has been reduced to 60mW by changing one of the mirrors in the zig-zag to a 33% beam splitter.  Though about 450mW is incident on the beam splitter the reflected beam is only about 55mW since the mirror reflectance is specified for P polarised light incident at 45deg while ours is S-polarised incident at less than 45deg.   The light transmitted through the beam splitter has been blocked by a beam dump.

 

 

  3800   Wed Oct 27 21:33:42 2010 SureshUpdateLockingGreen from the far end re-obtained

 

The mirror which was moved during the mode matching of PSL light to the MC (ref elog #3791) has been repositioned.  We once again have the green light from the NPRO on the X (south) arm available on the PSL table. 

This light was supposed to be collimated by the two plano convex lenses (f=200mm and f=50mm  ref to elog #3771) but it was converging.  So I moved the f=50mm lens backwards to make the beam collimated.  I checked the beam collimation by introducing an Al coated mirror infront of th PD and diverting the beam temporarily in a free direction.  I could then check the collinearity and collimation of both the green beams over a meter.  After alignment the mirror was removed and the light is now incident on the PD once again.  We can now proceed to look for green beats.

The power from the PSL NPRO was attenuated for the MC locking work of yesterday.  It has now been increased to the maximum by rotating the Half Wave Plate (HWP).  The power after the PSL is now about 450mW  (500mW - 10% picked off for the doubling).

 

 

 

  3799   Wed Oct 27 17:06:48 2010 josephbUpdateCDSMoved c1iscey chassis and host interface board to c1ioo

Problem:

Need a working IO chassis connected to c1ioo in order to bring the MC_L into the digital realm, and then via RFM transmit to the c1sus machine.

Attempted Solution:

Move the c1iscey IO chassis to c1ioo while the c1ioo chassis is at downs.

The c1iscey chassis however doesn't seem to be talking to the c1ioo computer.  I tried changing the host interface card on the c1ioo chassis.  I took out One Stop Systems HIB2-x4-H interface card with serial number 26638 from the c1ioo computer and put in the One Stop Systems HIB2-x4-H with serial number 35242 in from c1iscey into c1ioo.  Still didn't work.

All the lights are red on the interface card on the actual chassis and its cooling fan isn't spinning. 

Using dmesg on c1ioo shows that it does not see any of the ADC/DAC/BO cards.

Status:

I'm going to  wait until tomorrow morning when Rolf gets a chance to look at the c1ioo chassis over at Downs to determine the next step.  If we fix the c1ioo chassis, I'm move the c1iscey chassis and its host interface board back to the end.

  3798   Wed Oct 27 16:15:35 2010 SureshUpdateSUSRe-glued magnet to the PRM

Thanh and I re-glued the magnet to the PRM following the procedure outlined by Jenne

The PRM in the gluing fixture has been placed in the little foil house and left to cure for a day.

If all goes well the balancing the PRM will be done tomorrow.

 

 

  3797   Wed Oct 27 15:38:19 2010 josephb, yutaUpdateCDSIO chassis with bad timing was taken back to Downs

Problem:

The front end timing was not working properly for 2 of the IO chassis.  They were not being synced to the 1 PPS signal. 

This prevented the use of RFM for communication between front ends because time stamps on the transmitted data did not match the cycle on the receiving machine.

Action:

We took one of the incorrectly working chassis over to Downs.  Rolf said he would take a look at it tomorrow morning.

Joe will be going over tomorrow morning to talk with Rolf and see what needs to be done to fix it.

 

  3796   Wed Oct 27 12:32:53 2010 josephbUpdateCDSfb rebooted to try and fix testpoints

Problem:

Test points were unavailable last night, even after reboots of c1sus and even restarting the daqd process on the frame builder.

Cause:

Its unclear at this time.  My guess is flaky fb and mx_stream codes.  At the moment, the daqd often requires several restarts as it segfaults within a minute or two of restarting it.

What we did (aka treating the symptoms):

We rebooted the frame builder machine.  I also added the daqd and nds processes to the inittab.  Now when these die, they will automatically be restarted.

Steps to add to the inittab on fb

0) If not on fb, ssh -X fb

1) cd /etc/

2) sudo vi inittab or sudo emacs init

3) Add a line like: id:runlevels:action:process

The id is a unqiue 2-4 letter and number identifier for the process

Run levels is the run level of linux that it will start at. 345 will cover the normal cases

action is what to do with the process. Respawn makes it run at startup and also restarts it everytime it dies.

process is the command you want to run

See "man inittab" for more details

In this case we added

daq:345:respawn:/opt/rtcds/caltech/c1/target/fb/daqd -c /opt/rtcds/caltech/c1/target/fb/daqdrc > /opt/rtcds/caltech/c1/target/fb/daqd.log


nds:345:respawn:/opt/rtcds/caltech/c1/target/fb/nds pipe > /opt/rtcds/caltech/c1/target/fb/nds.log

4) Save.

5) Run "sudo /sbin/telinit q".  This forces init to rexamine the inittab file

daqd and nds will now automatically restart when they die.

Continuing issues:

When the frame builder dies, the mx_stream processes on the front ends die as well.  These need to be restarted manually at the moment by using "sudo /etc/restart_streams" while on c1sus.

The framebuilder code shouldn't be this flaky.

  3795   Wed Oct 27 11:52:45 2010 josephbUpdateelogElog needed to be restarted

I had to restart the elog on Nodus because it was no longer responding.

  3794   Wed Oct 27 11:34:40 2010 josephbUpdateCDSModified rc.local for front end machines

What was the problem:

On reboot, c1sus was in a strange state.  The epics IOC was running, along with tpman, but there were no loaded front ends.

People had to manually run sudo insmod c1SYSfe.ko.

What was the cause:

Awhile back Alex had commented out the line in the rc.local file which actually loaded the front end modules (i.e. c1x02fe.ko, c1susfe.ko, etc).  This was while debugging.  He never put it back.

What was the solution:

I uncommented the following line in /diskless/root/etc/rc.local file on the fb machine:

/opt/rtcds/caltech/c1/target/$i/scripts/startupC1rt

Now on reboot, the c1sus machine should start up all the front ends listed in the rtsystab file, located in /diskless/root/etc/ on the fb machine.

 

  3793   Wed Oct 27 10:53:03 2010 josephbConfigurationComputersWhy doesn't DTT work?!?

Test points for the SUS channels should be there.  They have been working previously this week.  Possibly break down points include awgtpman, mx_streams, and the fb itself.  I'll look into that.

As far as other fast channels, there are no other fast front ends running than the suspensions ones we have.  Until additional channels get connected to the front ends and the models updated, those are the channels we have available.  However I am working on getting c1ioo up and running, and we can try connecting in some PEM channels today to the c1sus front end's 4th ADC.

 

Edit:

I tried starting a fresh instance of the frame builder, but when I brought the old copy down, it left a pair of zombie or dead mx_stream processes running on c1sus . Basically c1mcs and c1rms were still running, while c1x02 and c1sus came down.  I tried to kill the processes but this caused the c1sus machine to crash.  In the past I've killed left over mx_stream processes running after the frame builder has gone down, but I've never seen them crash the computer.  I'm unsure why this happened since we haven't done any updates of the code, just updated models and daq configuration files.

Quote:

DTT has only SUS and "X02" channels under C1 in the drop down channel selection menu.  Basically, we can't measure any fast channels with DTT.  I keep getting the error: "Unable to select testpoints."  Sadface.

Similar things are true for DataViewer.  The same limited number of fast channels, and no data found:

Server error 13: no data found
datasrv: DataWriteRealtime failed: daq_send: Illegal seek

Is this a framebuilder problem?  Is this something that the CDS team has on the to-do list?

 

  3792   Wed Oct 27 09:02:29 2010 steveUpdateElectronicsHP4195A is NOT fixed

www.avalontest.com has fixed the 25MHz oscillation.  Contact: Jim Burnham 760-536-0191

Actually, NOT FIXED.

  3791   Wed Oct 27 02:26:15 2010 kiwamuUpdateIOOMC locked

(Rana, Koji, Suresh, Yuta, Thanh, Kiwamu)

 MC was locked successfully !

 Instead of feeding back the signal to the MC length we just injected it to the NPRO pzt with a high voltage (HV) amplifier.

So now we can move on to an in-vac work which needs the main beam to align the stuff.

 


(mode matching to MC)

Suresh and Thanh (a visitor from ANU) improved the mode matching to the MC. 

As written in the entry #3779  the beam after the mode matching lenses were diverging.

It is supposed to converge from 1.7mm radius at the last lens to 1.6 mm radius at the middle point of MC1 and MC2.

They slided the last lens toward the MC to make it more collimated and roughly measured the beam size using a sensor card.

As a fine tuning, they looked at some higher order modes showing up in the MC2 camera, and tried reducing the higher order modes by slightly sliding the last lens.

(assuming the lens position doesn't so much change the alignment)

During the work we removed a steering mirror for green locking because it was on the way of the lens slider.

 

- - measured optics' distances - - 

25.5 cm  from 1st lens to the front surface of the EOM

5.5 cm  length of the EOM

24.5 cm from the front surface of the EOM to the 2nd lens (concave)

15.5 cm  from the 2nd lens to the 1st steering mirror in the zig-zag path

20.5 cm  from the steering mirror to the last lens

 

(preparation for locking) 

Rana, Yuta and Koji prepared an old instant amplifier which can produce +/-13V output instead of usual SR560s.

We added an offset (~5V) on the signal to make it within 0-10V which is the input range of the HV amplifier.

If we take SR560, it's probably not sufficiently wide range because they can handle handle only about +/-4V.

 

We strung a cable from Marconi via the RF stabilizer to the wideband EOM in order to drive the EOM at 24.5MHz.

SInce the EOM doesn't have 50 Ohm input impedance we had to put a 50 Ohm load just before the EOM in order to drive it efficiently.

From a medm screen we set the driving RF amplitude slider (C1:IOO_MCRF_AMPADJ) to 0.0, which provides the maximum RF power on it.

 

(locking mode cleaner)

At first we unlocked the PMC to see an offset in the error signal without any lights on the MC_REFL PD.

Then we adjusted the offset to zero on the MC servo screen.

At the beginning of the locking the PMC was not stable for some reasons during the MC was locked.

But after increasing the laser power to the MC twice bigger, it looks like the PMC and the MC are quite stable.

 

 

  3790   Tue Oct 26 22:57:37 2010 JenneConfigurationComputersWhy doesn't DTT work?!?

DTT has only SUS and "X02" channels under C1 in the drop down channel selection menu.  Basically, we can't measure any fast channels with DTT.  I keep getting the error: "Unable to select testpoints."  Sadface.

Similar things are true for DataViewer.  The same limited number of fast channels, and no data found:

Server error 13: no data found
datasrv: DataWriteRealtime failed: daq_send: Illegal seek

Is this a framebuilder problem?  Is this something that the CDS team has on the to-do list?

  3789   Tue Oct 26 21:27:02 2010 JenneUpdateCDSfixed OPLEV stuff and MCL filters

Since MC2_TRANS is, in fact, MC2 Transmission, and not an oplev at all (it's not red, and it's not a lever, although it does use a QPD), I propose that the name be changed to something sensical, since calling it an oplev is completely non-sensical.  The name change should happen sooner rather than later, to avoid confusion.

Quote:

(Joe, Yuta)

Background:

 Today, we fixed OPLEV stuff, MCL filters, and time stamps.

What we did:
 5. Checked OPLEV for MC2 by pointing a laser pointer to QPD. (For MC2, OPLEV is just a transmission beam position monitor)
  Each quadrant looked like they are connected to the right channel numbers.

 

  3788   Tue Oct 26 17:31:17 2010 yutaUpdateCDSfixed OPLEV stuff and MCL filters

(Joe, Yuta)

Background:

 We are currently working on getting rid of "white stuff" in MEDM screens.
 Today, we fixed OPLEV stuff, MCL filters, and time stamps.

What we did:
 1. Plugged in OPLEV cables to ADC2. (See this wiki page for wiring)

 2. Connected ADC2 and OPLEV in Simulink model and fixed MEDM screens for OPLEVs (Attached #1).

 3. Put MCL filters for BS,ITMX,ITMY,PRM,SRM.
  They don't need them, but just for getting rid of "white stuff."
  They are connected to the ground, so the outputs are always 0.

 4. Fixed "TIME_STRING"s in MEDM screens so that they show current time correctly.
  You only need to put text monitor with channel "C1:FEC-DCU_NODE_ID_TIME_STRING" into master files(DEFAULTNAME things) and run generate_master_screens.py.
  It will automatically sets DCU ID correctly!! (Great work, Joe!)
  See this wiki page for more info on making MEDM screens.

 5. Checked OPLEV for MC2 by pointing a laser pointer to QPD. (For MC2, OPLEV is just a transmission beam position monitor)
  Each quadrant looked like they are connected to the right channel numbers.

Plan:
 - figure out what C1:SUS-NAME_MODE_SW1 does and fix
 - fix Whitening, Dewhitening ON/OFF button in main MEDM screens, so that they switch every channels' filters
 - make a new screen for MC (like the old one C1IOO_ModeCleaner.adl)
 - create a new mark for new MEDM screens

Attachment 1: optlevMC2.png
optlevMC2.png
  3787   Tue Oct 26 16:02:55 2010 yutaHowToCDSthings to do after making a new Simulink model

(Joe,Yuta)

Things to do after making a new Simulink model.

1. ssh c1sus, go to /opt/rtcds/caltech/c1/core/advLigoRTS/ and run
 bash ./makestuff.sh c1SYS

 makestuff.sh does;

  make uninstall-daq-$*
  make clean-$*
  make $*
  make install-$*
  make install-daq-$*
  make install-screens-$*
  sed -i 's/RO \(.*SW[12]R.*\)/\1/' /opt/rtcds/caltech/c1/target/$*/$*epics/autoBurt.req


 If you don't need to re-install DAQs or screens, just run line 2,3,4, and 7 and go to step #6.

2. ssh c1sus, go to /opt/rtcds/caltech/c1/scripts/ and run
 sudo ./startc1SYS

 For now, you have to put "1" in "BURT Restore" in GDS screens with in 5-10 secs.

3. Now the DAQ channel numbers are changed. So, go to /cvs/cds/rtcds/caltech/c1/chans/daq/ and run
 ./activateDAQ.py

 activateDAQ will activate the following DAQ channels for every optics with data rate 2048Hz;
  ULSEN_OUT
  URSEN_OUT
  LRSEN_OUT
  LLSEN_OUT
  SDSEN_OUT
  SUSPOS_IN1
  SUSPIT_IN1
  SUSYAW_IN1

4. Restart fb. See this wiki page.
 Basically you what have to do is kill and restart daqd in fb and restart mx_streams in c1sus.

5. DONE! Burt restore if you want.


6. If you don't need to re-install DAQs or screens;
  a. Go to /opt/rtcds/caltech/c1/target/c1SYS/c1SYSepics and run
    sudo rmmod c1SYSfe

  b. Go to /opt/rtcds/caltech/c1/target/c1SYS/bin/ and run
    sudo insmod c1SYSfe.ko

  3786   Tue Oct 26 15:57:10 2010 JenneUpdateSUSOne magnet broken, reglued

[Jenne, Suresh, Thanh (Bram's Grad Student)]

When we removed the grippers from the magnets on the PRM, one of the face magnets broke off.  This time, the dumbbell remained glued to the optic, while the magnet came off.  (Usually the magnet and dumbbell will stay attached, and both come off together).  I had 3 spare magnet-dumbbells, but only one of them was the correct polarization.  The strength of the spare magnet was ~128 Gauss, while the other magnets glued to the PRM are all ~180 Gauss.  We considered this too large a discrepancy, and so elected to reuse the same magnet as before. 

We removed the dumbbell from the optic using acetone.  After the epoxy was gently removed, we drag wiped the AR face of the optic (Acetone followed by Iso, as usual), being careful to keep all the solvent away from all the other glue joints.  We cleaned off the magnet with acetone (it didn't really have any glue stuck on it...most of the glue was stuck on the dumbbell), and epoxied it to a new dumbbell. 

The PRM, as well as the magnet-dumbbell gluing fixture are in the little foil house, waiting for tomorrow's activities.  Tomorrow we will re-glue this magnet to the optic, and Thursday we will balance the optic.  

This still leaves us right on schedule for giving the PRM to Bob on Friday at lunchtime, so it can bake over the weekend.

  3785   Tue Oct 26 12:45:20 2010 yutaHowToCDSmaking a new master file for medm screens

(Joe, Yuta)

 In elog #3708, we showed how to edit all the similar medm screens easiliy.
 But if there are no master files you want to edit in /opt/rtcds/caltech/c1/medm/master/ directory, you can make one.
 Say, you want to edit all the similar medm screens named C1SUS_NAME_XXX.adl.

1. Copy one of C1SUS_NAME_XXX.adl to /opt/rtcds/caltech/c1/medm/master/ directory as C1SUS_DEFAULTNAME_XXX.adl.

2. Go to that directory and
  sed -i 's/NAME/DEFAULTNAME/g' C1SUS_DEFUALTNAME_XXX.adl

3. Add "_XXX.adl" to file_array list in generate_master_screens.py

4. Edit C1SUS_DEFUALTNAME_XXX.adl

5. Run ./generate_master_screens.py

That's all!!

  3784   Tue Oct 26 10:50:08 2010 KojiConfigurationelogELOG 2.8.0 -> ELOG 2.7.5 -> ELOG 2.8.0

ELOG restarted with 2.8.0 again.

- moved elog-2.8.0/script dir to elog-2.8.0/script.orig

- copied elog-2.7.5/script to elog-2.8.0/script

- /cvs/cds/caltech/elog/start-elog.csh reconfigured to launch 2.8.0

- /cvs/cds/caltech/elog/elog is linked to ./elog-2.8.0

- logbooks on 25th and 26th were copied from 2.7.5 to 2.8.0.

 

  3783   Tue Oct 26 07:02:05 2010 AlbertoConfigurationelogELOG 2.8.0 -> ELOG 2.7.5

Quote:

ELOG reverted to 2.7.5 due to editing difficulties

- /cvs/cds/caltech/elog/start-elog.csh reconfigured to launch 2.7.5

- /cvs/cds/caltech/elog/elog is linked to ./elog-2.7.5

- logbook dir of 2.8.0 was copied in the dir of 2.7.5. The old and obsolete 2.7.5 was discarded.

I think I had the same problem when I switched to 2.75 from 2.65.

Then the problem was FCKeditor.

We should try the solution I put in the elog page of the wiki.

 

  3782   Tue Oct 26 01:53:21 2010 Joonho LeeUpdateElectronicsFuction Generator removed.

Today I worked on how to measure cable impedance directly.

In order to measure the impedance in RF range, I used a function generator which could generate 50MHz signal and was initially connected to the table on the right of the decks.

The reason I am checking the cables is for replacing the cables with impedance of 50 or 52 ohm by those with impedance of 75 ohm.

After I figures out which cable has not proper impedance, I will make new cables and substitute them in order to match the impedance, which would lead to better VIDEO signal.

 

To test the VIDEO cables, I need a function generator generating signal of frequency 50 MHz.

In the deck on the right of PSL table, there was only one such generator which was connected to the table on the right of the deck.

Therefore, I disconnected it from the cable and took it to the control room to use it because Rana said it was not used.

Then, I tired to find on how to measure the impedance of cable directly but I did not finish yet.

When I finished today works, I put the generator back to the deck but I did not connect to the previous cable which was initially connected to the generator.

 

Next time, I will finish the practical method of measuring the cable impedance then I will measure the cables with unknown impedance.

Any suggestion would be appreciated.

  3781   Tue Oct 26 00:45:08 2010 ranaUpdatePSLNew NPRO Diagnostic Wiring

Untitled.png

I copy section 6.2 into here to share with you all what the diagnostic capabilities of the new NPRO are. Its not a lot.

We'll need to record a sample of the NPRO output beam on a regular photodiode in order to get a real power monitor. My plan is to use a regular 25-pin Dsub and run it fron the NPRO controller over to the PSL rack and hijack the old MOPA monitoring channels (3113 and 3123 ADCs).

  3780   Mon Oct 25 23:59:37 2010 KojiConfigurationelogELOG 2.8.0 -> ELOG 2.7.5

ELOG reverted to 2.7.5 due to editing difficulties

- /cvs/cds/caltech/elog/start-elog.csh reconfigured to launch 2.7.5

- /cvs/cds/caltech/elog/elog is linked to ./elog-2.7.5

- logbook dir of 2.8.0 was copied in the dir of 2.7.5. The old and obsolete 2.7.5 was discarded.

  3779   Mon Oct 25 23:10:06 2010 KojiUpdateIOOMC is now flashing

[Suresh / Koji]

The MC mirrors are aligned. Now the flashing of the resonances are visible on the MC2 CCD

although the modematching seemed pretty poor.


- The incident power was adjusted to be ~20mW by rotating HWP after the laser source.
The power before the window of the chamber was ~450mW. Where are those missing 1.5W?

- We checked the spot on the last two steering mirrors and the incident beam on MC1.
The beam was too much off from the center of the 1st steering mirror. It was also hitting 1cm north of the MC1.
We adjusted the steering mirrors such that the incident and reflected beams are symmetrically visible at the MC1 tower.

- The MC mirrors are aligned. We first tried to use only MC2 and MC3. And then we used MC1 too as the spot on the MC2 was too high.

- We saw some TEM00 flashes but with many other modes flashing. We checked the beam diameter on the PSL table and on the MC REFL.
The latter one looked twice large as the former one. We concluded the beam is diverging.

- We closed the tank and decided to work on the mode matching tomorrow.

  3778   Mon Oct 25 20:20:59 2010 yuta, JoeUpdateCDSDAQ channel number etc
Summary:
We split the old SDSEN filters to SDSEN, SDSIDE, SDCOIL last week.
Along with this change, the TP channel number changed unfortunately.
So, we fixed them.
Also, we made FM9 do the output filter analog/digital switching.

What we did:
1. Changed the Simulink logic so that FM9 do the output filter switching, and checked the logic by probing
MAX333A for SDCOILs.

2. After making a new Simulink model and rebuilding, run the following incantation to burt restore filter
settings in /opt/rtcds/caltech/c1/target/c1SYS/c1SYSepics/ (See elog #3706)
sed -i 's/RO \(.*SW[12]R.*\)/\1/' autoBurt.req

3. DAQ channel numbers are listed in C1SYS.ini files in /cvs/cds/rtcds/caltech/c1/chans/daq/.
Channels with # signs are not activated. So, we changed, for example,

#[C1:SUS-MC1_LLSEN_IN1_DAQ]
#acquire=0
#datarate=16384
#datatype=4
#chnnum=10208

to

[C1:SUS-MC1_LLSEN_IN1_DAQ]
acquire=1
datarate=2048
datatype=4
chnnum=10208

for all channels we want to look at.

4. Restarted fb.

Plan:
- measure TFs and see if input and output filter switchings are working correctly
- make a switch that does all filter switching for all 5 OSEMS or 5 COILS
- put optical lever stuff
- fix offset sliders and offset switch
- put F2A filters in TO_COIL matrix (see elog #3769)
- make a nice graphical screen for MCs (like /cvs/cds/caltech/medm/c1/ioo/C1IOO_ModeCleaner.adl)
- write a script that activates DAQ we need
- make a plan

* SYS=SUS, RMS, MCS, etc
  3777   Mon Oct 25 11:39:06 2010 JenneUpdateSUSMagnets glued to PRM
This morning I glued the magnets to the PRM. Now we wait, and tomorrow afternoon (at the earliest), Suresh and I can balance the PRM.
Attachment 1: StatusTable.png
StatusTable.png
  3776   Mon Oct 25 02:25:21 2010 KojiUpdatePSLQuarter Wave Plate Measurements

Q1. Suppose the laser beam has a certain (i.e. arbitrary) polarization state but contains only TEM00. Also suppose the PSB is perfect (reflect all S and transmit all P). What results do you expect from your expereiment?

Q2. Suppose the above condition but the PBS is not perfect (i.e. reflects most of S but also small leakage of P to the reflection port.) How are the expected results modified?

Q3. In reality, the laser may also contain some thing dirty (e.g. deporarization in the laser Xtal, higher order modes in a certain polarization but different from the TEM00's one, etc). What actually is the cause of 170mW rejection from the PBS? Can we improve the transmitted power through the PBS?

Q4. Why is the visibility for the lambda/4 with 330deg better than the one with 326deg? Yes, as I already explained to Kevin, I suppose it was caused by the lack of the data points in the wider angle ranges.

Quote:

I measured the reflected power from the PBS as a function of half wave plate rotation for five different quarter wave plate rotations.

The optimum angles that minimize the reflected power are 330° for the quarter wave plate and 268° for the half wave plate.

The following data was taken with 2.102 A laser current and 32.25° C crystal temperature.

For each of five quarter wave plate settings around the optimum value, I measured the reflected power from the PBS with an Ophir power meter. I measured the power as a function of half wave plate angle five times for each angle and averaged these values to calculate the mean and uncertainty for each of these angles. The Ophir started to drift when trying to measure relatively large amounts of power. (With approximately 1W reflected from the PBS, the power reading rapidly increased by several hundred mW.) So I could only take data near the minimum reflection accurately.

The data was fit to P = P0 + P1*sin^2(2pi/180*(t-t0)) with the angle t measured in degrees with the following results:

lambda/4 angle (°) t0 (°) P0 (mW) P1 (mW) chi^2/ndf V
318 261.56 ± 0.02 224.9 ± 0.5 2016 ± 5 0.98 0.900 ± 0.001
326 266.07 ± 0.01 178.5 ± 0.4 1998 ± 5 16.00 0.918 ± 0.001
330 268.00 ± 0.01 168.2 ± 0.3 2119 ± 5 1.33 0.926 ± 0.001
334 270.07 ± 0.02 174.5 ± 0.4 2083 ± 5 1.53 0.923 ± 0.001
342 273.49 ± 0.02 226.8 ± 0.5 1966 ± 5 1.41 0.897 ± 0.001

where V is the visibility V = 1- P_max/P_min. These fits are shown in attachment 1. We would like to understand better why we can only reduce the reflected light to ~150 mW. Ideally, we would have V = 1. I will redo these measurements with a different power meter that can measure up to 2 W and take data over a full period of the reflected power.

 

  3775   Mon Oct 25 02:23:47 2010 KojiConfigurationelogELOG 2.8.0
When I push the reply button, the raw html shows up in the edit window and have to use HTML to write the entry.
Does this happen only to me???


Quote:
I stopped the ELOG and restarted us on 2.8.0.

To make sure nothing got lost, I killed the old process, copied over the logbooks/, themes/, and elogd.cfg to the new 2.8.0/ directory before starting the new Daemon.

I encountered the same Administrator bug as Joe had before. I delete all the old Admin passwords to bypass the issue.

To restart the ELOGD on NODUS, you now type '/cvs/cds/caltech/elog/start-elog.csh'.
I also added ELOG to the man pages in /usr/local/man/ on nodus by putting the *.1 files in man1/ and the *.8 files into man8/.
  3774   Mon Oct 25 02:14:38 2010 KojiSummaryLockingsetup for green beat

- What is the actual photocurrent for the beam1 and beam2? We don't care how much power do you have before the BS, but care how much current do you have on the PD.

- How much is the visibility? There is mismatching of the beams. i.e. The beam diameter looked quite different. Also the beams are not TEM00 but have fringes probably comes from the TT mirrors. You maybe able to measure the visibility by the DC output, making the beat freq go through df=0 slowly.

- What is the measured gain of the RF amp? Does it include the voltage division by the output/input impedance?

---------------------

 The signal level of the observed peak was -48dBm.

However I was expecting it is like -28dBm with some ideal assumptions.

There may be a 20dB unknown loss which sounds big to me.

I was assuming the parameters are like:

           A = 0.39 [A/W]   (assuming 90% quantum efficiency at 532nm)

           Z = 240 [V/A]  

           P1 = 17 uW  (measured by Newport power meter)

           P2 = 30 uW (measured by Newport power meter)

           G_RF = 23 dB

  3773   Sun Oct 24 19:55:50 2010 kiwamuUpdateElectronicslonely RF amplifier on ITMX table
(Rana, Kiwamu)

Last Friday we found a lonely RF amplifier ZHL-3A on the ITMX table.
When we found him we were very sad because he's been setup unacceptably


For example, the signal input was disconnected although a 24V DC was still applied. So he has been making just a heat for a long time.
The power connector was a BNC style which is not official way.
The leg of a decoupling capacitor attached to the DC connector was apparently broken and etc,..

We salvaged him and then cleaned up those cables and the DC power supply.

We don't say like 'don't make a temporary setup', but please clean up them after finishing the work every time.
  3772   Sun Oct 24 19:23:41 2010 ranaConfigurationelogELOG 2.8.0
I stopped the ELOG and restarted us on 2.8.0.

To make sure nothing got lost, I killed the old process, copied over the logbooks/, themes/, and elogd.cfg to the new 2.8.0/ directory before starting the new Daemon.

I encountered the same Administrator bug as Joe had before. I delete all the old Admin passwords to bypass the issue.

To restart the ELOGD on NODUS, you now type '/cvs/cds/caltech/elog/start-elog.csh'.
I also added ELOG to the man pages in /usr/local/man/ on nodus by putting the *.1 files in man1/ and the *.8 files into man8/.
  3771   Sun Oct 24 18:06:35 2010 kiwamuSummaryLockingsetup for green beat

green_beat_note.png

(notes on signal level)

 The signal level of the observed peak was -48dBm.

However I was expecting it is like -28dBm with some ideal assumptions.

There may be a 20dB unknown loss which sounds big to me.

 

The expectation was calculated by using the following simple math.

                       SIGNAL = A x Z x G_RF x sqrt(P1 / 2) x sqrt (P2 / 2)

 where A is the responsibility of the PD and Z is the trans impedance gain. G_RF is a gain of the RF amplifier.

The laser powers of green beams, P1 and P2, are divided by 2 due to a beam splitter. 

I was assuming the parameters are like:

           A = 0.39 [A/W]   (assuming 90% quantum efficiency at 532nm)

           Z = 240 [V/A]  

           P1 = 17 uW  (measured by Newport power meter)

           P2 = 30 uW (measured by Newport power meter)

           G_RF = 23 dB

 

  3770   Sat Oct 23 14:42:01 2010 yutaSummaryCDSdamped MC suspensions

After replacing filters, MC suspensions damped  last night.

Further measurement next time.

Attachment 1: dam.png
dam.png
  3769   Sat Oct 23 03:36:05 2010 yutaUpdateCDSfixed filters for C1SUS, C1RMS, C1MCS

(Joe, Yuta)

Summary:

 This Monday, MC suspension damping got something wrong.
 We started to check filters and found that digital filters were wrong because of mis-conversion from old filter files to new files.
 We converted the file again, and with mutual understanding between Foton and us, we finally got correct filters(I hope!).

What we did:
 1. Merged filter files in old /cvs/cds/caltech/chans/ directory into C1SUS.txt(BS,ITMX,ITMY), C1RMS.txt(PRM,SRM), C1MCS.txt(MC123).

 2. Rebuilt the RT models in order to get a correct filter file header(they have list of filter modules).

 3. Concatenate the header with filter design part which we got from step1.

 4. Replaced 'N 2048' with 'N 16384'
   It replaces sampling rate of "XXSEN"s.

Basically, step1-4 was the same with what we did last time. We didn't changed the fitler coefficients, so Foton somehow changed the original filter design.
So, this time, we

 5. Deleted coefficients like we did on Tuesday (see elog #3774).

But Foton couldn't read the file correctly this time. Foton seemed to be unbeatable.
Even if we replaced the sampling rate, Foton kept saying 2048! (This is maybe because Foton's default value is 2048Hz. Everytime Foton notice some editting in the file, he destroys everything. He hates editting)
The problems were always associated with the sampling rate, so

 6. Got back to step4 and undo-ed the replacement.

 7. Foton could read it this time, so I changed the sampling rate one by one using Foton GUI.

 8. Checked filters using Foton's Bode Plot.(Not for all, but some that had problem before)

 9. Splitted SDSEN filters to SDSEN, SUSSIDE, and SDCOIL.

 10. Put some missing whitening filters, and 28HzELPs.
   BS, PRM, SRM didn't have any 28HzELP for SDCOIL.
   ITMX and ITMY SDCOIL had SimDW/InvDW which doesn't make sense(SIDEs don't have analog DW). So, I deleted and replaced with 28HzELP.

F2A issue:
 We failed in sending F2A filters to new filter files.
 These are a little bit complicated because TO_COIL_X_X filters were named ULPOS,URPOS etc before.
 Also, MC3 didn't have any F2As, so maybe we should but the same F2As as MC1/2.
 Note that every F2As are different, and TO_COIL matrix have UL,UR,LL,LR order(not same as INMATRIX).
 Also, SRM had f2pv instead of F2A!

Next work:
 - Check whole filters by actually measuring transfer function between SENs and COILs.
 - Damp MC suspentions, and lock MC.
 - Measure openloop TF and compare with the designed.


How do you read a Foton filter file:
 When you open up a Foton filter file, you see filters like this.

################################################################################
### modulename                                                               ###
################################################################################
# SAMPLING modulename samplingrate
# DESIGN   modulename n filterdesign
# DESIGN   modulename n filterdesign
###                                                                          ###
modulename   n xy z      v      w filtername                        gain    a1     a2    b1     b2
modulename   n xy z      v      w filtername                        gain    a1     a2    b1     b2
                                                                            a1     a2    b1     b2
n: filter number
 0 for FM1, 1 for FM2, ... , 9 for FM10

x: Input Switching setting
 1 Always On
 2 Zero History

y: Output Switching setting
 1 Immediately
 2 Ramp
 3 Input Crossing
 4 Zero Crossing

z: number of filters cascaded.

v: if y=2, (Ramp Time(sec))*(samplingrate)
   if y=3 or 4, Tolerance

w: (Timeout(sec))*(samplingrate)

 Note that v and w are changed when sampling rate is changed.

 Transfer function will be;
  H(1/z)=G*(1+b1/z+b2/z/z)/(1+a1/z+a2/z/z)
  z=exp(s/fs)

 where fs is the sampling frequency.

Reference:
 Kiwamu Izumi: "Notes about Digital Filters," http://tamago.mtk.nao.ac.jp/izumi/green/DigitalFilter.pdf

  3768   Sat Oct 23 02:25:49 2010 KevinUpdatePSLQuarter Wave Plate Measurements

Quote:

[Koji and Kevin]

We measured the reflection from the PBS as a function of half wave plate rotation for five different quarter wave plate rotations. Before the measurement we reduced the laser current to 1 A, locked the PMC, and recorded 1.1 V transmitted through the PMC. During the measurements, the beam was blocked after the faraday isolator. After the measurements, we again locked the PMC and recorded 1.2 V transmitted. The current is now 2.1 A and both the PMC and reference cavities are locked.

I will post the details of the measurement tomorrow.

I measured the reflected power from the PBS as a function of half wave plate rotation for five different quarter wave plate rotations.

The optimum angles that minimize the reflected power are 330° for the quarter wave plate and 268° for the half wave plate.

The following data was taken with 2.102 A laser current and 32.25° C crystal temperature.

For each of five quarter wave plate settings around the optimum value, I measured the reflected power from the PBS with an Ophir power meter. I measured the power as a function of half wave plate angle five times for each angle and averaged these values to calculate the mean and uncertainty for each of these angles. The Ophir started to drift when trying to measure relatively large amounts of power. (With approximately 1W reflected from the PBS, the power reading rapidly increased by several hundred mW.) So I could only take data near the minimum reflection accurately.

The data was fit to P = P0 + P1*sin^2(2pi/180*(t-t0)) with the angle t measured in degrees with the following results:

lambda/4 angle (°) t0 (°) P0 (mW) P1 (mW) chi^2/ndf V
318 261.56 ± 0.02 224.9 ± 0.5 2016 ± 5 0.98 0.900 ± 0.001
326 266.07 ± 0.01 178.5 ± 0.4 1998 ± 5 16.00 0.918 ± 0.001
330 268.00 ± 0.01 168.2 ± 0.3 2119 ± 5 1.33 0.926 ± 0.001
334 270.07 ± 0.02 174.5 ± 0.4 2083 ± 5 1.53 0.923 ± 0.001
342 273.49 ± 0.02 226.8 ± 0.5 1966 ± 5 1.41 0.897 ± 0.001

where V is the visibility V = 1- P_max/P_min. These fits are shown in attachment 1. We would like to understand better why we can only reduce the reflected light to ~150 mW. Ideally, we would have V = 1. I will redo these measurements with a different power meter that can measure up to 2 W and take data over a full period of the reflected power.

Attachment 1: fits.png
fits.png
  3767   Fri Oct 22 23:29:48 2010 yutaUpdateCDSfixed output filter switching

(Joe, Yuta)

Summary:
 This evening, we fixed output filter analog/digital switchings in Simulink logic, but they didn't actually switched filters.
 That was because what we thought BO0 was BO2 and BO2 was BO0.
 We re-wired, re-labeled the cables.
 We checked it by using a LED(see "test configuration" in this wiki page).
 Also, we checked the Simulink connection by probing MAX333A in SOS Dewhitening Board located at 1X4-8-11B, which switches SIDE coils for all optics.

Next work:

 - Currently, FM10 in UL/UR/LR/LL/SDCOIL filters switch analog/digital, but let's make FM9 do the switching.(See schematic in elog #3758)
 - Check all the logic by measuring transfer functions(maybe single frequency measurement is enough. we can use tdsdmd etc to automate).

  3766   Fri Oct 22 21:28:27 2010 AidanUpdateLockingGreen era: found a green beat note finally

Nice work!

Quote:

finally we found it !

 green_beatnote.jpg

 

  3765   Fri Oct 22 19:53:27 2010 yutaSummaryCDSconversion failure in digital filters

(Rana, Joe, Yuta)

We now understand that we never succeeded in converting old fiter files to new filter files.

For example, we just changed the sampling rate with coefficients remained and foton confused, or we forgot to rename some of the module names(ULSEN -> SUS_MC1_ULSEN) ......

This is why we sometimes damped and sometimes didn't, depending on filter switches. New filter files has been always wrong.

So, we started to convert them again.
We have to figure out how to convert the files that Foton accepts correctly.

  3764   Fri Oct 22 18:22:27 2010 AlbertoUpdateElectronicsEpic Takeover

Quote:

As the suspension work winds down (we'll be completely done once the ETMs arrive, are suspended, and then are placed in the chambers), I'm going to start working on the RF system. 

Step 1: Figure out what Alberto has been up to the last few months.

Step 2: Figure out what still needs doing.

Step 3: Complete all the items listed out in step 2.

Step 4: Make sure it all works.

Right now I'm just starting steps 1 & 2.  I've made myself a handy-dandy wiki checklist: RF Checklist.  Hopefully all of the bits and pieces that need doing will be put here, and then I can start checking them off. Suggestions and additions to the list are welcome.

 There's also a page dedicated to the progress in the PD upgrade process:

http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/RF_System/Upgraded_RF_Photodiodes

There you can find a pdf document with my notes on that.

  3763   Fri Oct 22 18:15:21 2010 kiwamuUpdateLockingGreen era: found a green beat note finally

finally we found it !

 green_beatnote.jpg

  3762   Fri Oct 22 16:59:21 2010 JenneUpdateElectronicsEpic Takeover

As the suspension work winds down (we'll be completely done once the ETMs arrive, are suspended, and then are placed in the chambers), I'm going to start working on the RF system. 

Step 1: Figure out what Alberto has been up to the last few months.

Step 2: Figure out what still needs doing.

Step 3: Complete all the items listed out in step 2.

Step 4: Make sure it all works.

Right now I'm just starting steps 1 & 2.  I've made myself a handy-dandy wiki checklist: RF Checklist.  Hopefully all of the bits and pieces that need doing will be put here, and then I can start checking them off. Suggestions and additions to the list are welcome.

  3761   Fri Oct 22 15:06:43 2010 JenneUpdateSUSArts and Crafts!

This afternoon I epoxied the guiderod and wire standoff to the new PRM.  I also epoxied the magnets that Suresh picked out to the dumbbell standoffs.  We'll let them all cure over the weekend, and then I'll glue the magnets to the optic on ~Monday.

Notes about the epoxy: 

Previously, we had been using the "AN-1" epoxy, which is gray, with a clear hardener.  Bob recommended we switch to "30-2", which is clear with clear, and has been chosen for use in aLIGO.  Both were vacuum approved, but the 30-2 has gone through ~2 months of testing at the OTF (Optics Test Facility?) over in Downs under vacuum, to check the level of outgassing (or really, non-outgassing).

The 30-2 is less viscous than the AN-1, and it takes less glue to do the same job, so we should keep that in mind when applying the epoxy.  When I put the glue next to the guiderod and standoff, it got wicked along the length of each rod, which is good.  I can't reach the whole length of the rod with my glue applicator because the fixture holding them in place blocks access, so the wicking is pretty handy.

 

I've also added the updated version of my Status Table for the suspensions.

Attachment 1: StatusTable.png
StatusTable.png
  3760   Fri Oct 22 03:37:56 2010 KevinUpdatePSLQuarter Wave Plate Measurements

[Koji and Kevin]

We measured the reflection from the PBS as a function of half wave plate rotation for five different quarter wave plate rotations. Before the measurement we reduced the laser current to 1 A, locked the PMC, and recorded 1.1 V transmitted through the PMC. During the measurements, the beam was blocked after the faraday isolator. After the measurements, we again locked the PMC and recorded 1.2 V transmitted. The current is now 2.1 A and both the PMC and reference cavities are locked.

I will post the details of the measurement tomorrow.

  3759   Fri Oct 22 01:23:13 2010 KojiUpdatePSLFound the beat at 1064nm

[Koji / Suresh]

We worked on the 1064 beating a bit more.

- First of all, FSS and FSS SLOW servo were disabled not to have variating Xtal temp for the PSL.

- The PSL Xtal temp (T_PSL) was scanned from 22deg-45deg while we search the Xtal temp (T_Xend) for the Xend laser to have the beat freq well low (f<30MHz).
The pumping current for each laser was I_PSL = 2.101 [A] and I_Xend = 2.000 [A]

For a certain T_PSL, we found multiple T_Xend because the freq of the laser is not a monotonic function of the Xtal temperature. (see the innolight manual).

T_Xend to give us the beating was categorized in the three sets as shown in the figure. The set on "curve2" is the steadiest one. (Use this!)
The trends were quite linear but the slope was slightly off from the unity.

- T_PSL was scanned to see the trend of the PMC output.

The PMC was sometimes locked to the mode with lower transmission (V_PMCT ~ 3.0V).
When T_PSL ~ 31deg we consistently locked the PMC at higer transmission (V_PMCT ~ 5.3V).

At the moment we decided the operating point of T_PSL = 32.25 deg, V_PMCT = 5.34, where we found the beat at T_Xend=38.28deg.

- We cleaned up the PSL table more than how it was. Returned the tools to their original places.
The X-end laser was shut down and was left on the PSL table.

NEXT:
Kiwamu can move the X-end laser to the Xend and realign it.
Then we should be able to see the green beating quite easily.

Attachment 1: 101021_beat.pdf
101021_beat.pdf
  3758   Fri Oct 22 00:59:01 2010 yutaUpdateCDSnew MEDM screens, new filter banks(whitening and dewhitening related)

(Joe, Yuta)

Summary:
 No more discrimination for SIDE!
 We had all SIDE filters in SDSEN, so we splitted into SDSEN, SUSSIDE, and SDCOIL just like other DOFs. (Not SIDECOIL. SDCOIL from now on!)
 Also, there was a confusion in output filters, so we fixed the filter bank(dewhitening and 28HzELP).
 More than that, I found that "13HzCheby" in SUSPOS, SUSPIT, SUSYAW was actually doing ~1.6Hz Chebyshev, so I fixed it. (No wonder we had no damping when turing "13HzCheby" on!)
 Our new MEDM screen is Attachment #1.

Input filter design:
 Every optics, every OSEM has same analog whitening filter.
 So;

digital analog
30,100:3 30,100:3
ON OFF
OFF ON

Output filter design:
 For every SIDE coils and MC1, MC3 coils, they have analog 28Hz elliptic LPFs.
 So;
digital analog
28HzELP 28HzELP
ON OFF
OFF ON

 For MC2(and other main optics) UL/UR/LR/LL coils, they have analog dewhitening filters.
 So;
digital analog
InvDW SimDW DW
ON ON OFF
ON OFF ON

 InvDW and SimDW were in FM10 and FM9, but I swapped them to make it more consistent.
 Now, FM10 switches analog/digital output filiters for every optics.

  Let's put 28HzELP in FM9 so that FM9 switches analog/digital for every optics.

Quote:

Schematic:
 - whitening
   MC1 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC2 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
   MC3 5 PD outputs -> SUS PD Whitening Board(D000210) -> ... (digital WF=3,100:3)
      D000210 has switches for bypassing analog WT(3,100:3). HIGH to bypass.
 - dewhitening
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,3 UL/UR/LR/LL coils
  (-) ... -> SOS Dewhitening Board(D000316) -> MC1,2,3 SIDE coils
  (SimDW)(InvDW) ... -> LSC Anti-imaging Board(D000186) -> Universal Dewhitening Board(D000183) -> MC2 UL/UR/LR/LL coils
     D000316 has switches for bypassing 28Hz elliptic LPF. HIGH to bypass.
     D000186 is 7570Hz elliptic LPFs.
     D000183 has switches for bypassing dewhitening filter. HIGH to bypass.
 See this wiki page for more comprehensive setup.

Next work:
 - Fix Simulink connection and logic so that analog/digital switching go correctly.
 - Check if the switching are correct for all channels by measuring transfer functions.
 - Currently, we are focusing on MCs, but we have to do the same fixing for other optics, too.
 - More than setting the right filter TF, there are switching settings like Zero History, Ramp, Zero Crossing...... We should investigate that.
 - Actually, TF of analog DW in D000183 doesn't look like the same with digital SimDW. Why??

Attachment 1: HappySideMEDM.png
HappySideMEDM.png
  3757   Thu Oct 21 21:35:16 2010 yutaUpdateCDSfixed whitening filter switches

(Joe, Yuta)

Summary:
 We found some mistakes in whitening filter switches yesterday(see elog #3748).
 One was the mis-connection between LL/LR and the second was the bit invert.
 So, we fixed it and checked that they are now switching correct by probing MAX333A.

Why we made mistake:
1. LL/LR mis-connection
  Simply, Simulink model was wrong.
  It occurred because UL/UR/LR/LL order is currently arbitrary and confusing.
  For now, we just swap the connection(Attached #1).
  In the future, we should fix all orders corresponding to the actual wiring order(UL > LL > UR > LR).

2. Bit invert
  It occurred from counter-intuitive output of Contec 32 BO(See Attachment #2(taken from this wiki page)).
  If "1," current goes from A to B, so MAX333A input will be 0V, which means analog WF is on.
  If "0," no current, so MAX333A input will be +15V, which means analog WF is off(bypassed).

Next work:
 - There are no digital WF in SDSEN inputs. So, we have to put them.
 - Dewhitening filters are totally messed up now. Switches are wrong, some optics have digital DW but some doesn't, some have 28Hz elliptic LPF.......
     We have to grand unify them.

Attachment 1: LLLR.png
LLLR.png
Attachment 2: Contec23BO.png
Contec23BO.png
ELOG V3.1.3-